# MuLan: Adapting Multilingual Diffusion Models for Hundreds of Languages with Negligible Cost

### v1

@@ Abstract
In this work, we explore a cost-effective framework for multilingual image generation. We find that, unlike models tuned on high-quality images with multilingual annotations, leveraging text encoders pre-trained on widely available, noisy Internet image-text pairs significantly enhances data efficiency in text-to-image (T2I) generation across multiple languages. Based on this insight, we introduce MuLan, Multi-Language adapter,a lightweight language adapter with fewer than 20M parameters, trained alongside a frozen text encoder and image diffusion model. Compared to previous multilingual T2I models, this framework offers: (1) Cost efficiency. Using readily accessible English data and off-the-shelf multilingual text encoders minimizes the training cost; (2) High performance. Achieving comparable generation capabilities in over 110 languages with CLIP similarity scores nearly matching those in English (38.61 for English vs. 37.61 for other languages); and (3) Broad applicability. Seamlessly integrating with compatible community tools like LoRA, LCM, ControlNet, and IP-Adapter, expanding its potential use cases.
* equal contribution; ✉ corresponding author: wangwenhai@pjlab.org.cn.
@@ 1 Introduction
Recent diffusion models [11, 18, 19, 40, 44, 48] for content generation have attained stunning advancements in terms of both aesthetic quality and text-content alignment. However, these models still face substantial limitations in multilingual support. For instance, one of the most popular image-generation models, Stable Diffusion [33], and its successors [28, 11] only supports English and a few Latin-based languages. This language barrier restricts the model’s performance in multilingual contexts and hinders its applicability worldwide across diverse cultural and linguistic backgrounds.
@@
The community primarily adopts two approaches to achieve multilingual text-to-image (T2I) generation. (1) The first is translation-based methods where input content is temporarily translated into English before generating the image. However, this approach often leads to inference delays, translation errors, and notable issues when handling slang or culturally nuanced content. (2) The other approach is native multilingual T2I models [48, 29, 1, 39, 19, 40], which is trained directly on high-quality images captioned in the target language. While this native approach improves image generation quality for non-English languages, it relies on extensive, carefully curated image-generation data in the target language, making it data- and resource-intensive. As a result, exploring a more efficient and generalizable approach to achieve strong multilingual generation capabilities remains challenging.
@@
On the other hand, thanks to the development of computational power and dataset scale, many existing language models have achieved strong multilingual capabilities through training on large-scale internet data. For example, models trained on text data (e.g., BERT [8], the GPT series [2, 25], and LLaMA [42, 9]) and those trained on image-text pairs (e.g., CLIP [29], ALIGN [15], and InternVL [6]) demonstrate outstanding performance in multilingual understanding. Since the multilingual capability of T2I generation models is closely tied to their text encoders, it becomes essential to explore how these powerful multilingual text encoders can be leveraged to enable existing generative models to achieve multilingual capabilities more efficiently and effectively.
@@
In this work, we deeply explore the application of multilingual semantic alignment in image generation from the perspective of language and image-text alignment. We also reveal that text encoders trained on large-scale multilingual image-text datasets with noisy data demonstrate remarkable data efficiency in multilingual image generation. Based on this insight, we introduce MuLan, a lightweight Multi-Language adapter that connects text encoders to diffusion models with low-cost training, facilitating native support for hundreds of languages. Specifically, we employ a plug-and-play language adapter with fewer than 20 million parameters to bridge a frozen multilingual text encoder with a frozen diffusion model. This setup exhibits outstanding zero-shot performance on multilingual T2I generation tasks, even relying on minimal English training data. Additionally, the framework is highly compatible with widely used community models and tools, such as LoRA [13], LCM [23], ControlNet [49], and IP-Adapter [47], demonstrating excellent flexibility, as shown in Figure 1.
@@
Our model offers advantages in terms of training cost, multilingual performance, and adaptability compared to previous works. Specifically, MuLan leverages large-scale pre-trained multilingual text encoders, reducing the need for extensive multilingual datasets. As a result, the model can efficiently adapt to over 110 languages using only a small amount of English training data, achieving generation quality comparable to English (e.g., CLIP similarity scores: 38.61 for English, average 37.61 for other languages). Additionally, as shown in Figure 4, MuLan ’s plug-and-play architecture seamlessly integrates with existing model architectures and frameworks, greatly enhancing its compatibility with various community-driven tools and models.
@@
(1) We investigate different possible ways to equip monolingual text-to-image models with multilingual ability, among which we demonstrate that using a text encoder properly trained from noisy web-scale data (such as InternVL [6]) is of great data efficiency.
@@
(3) We demonstrate our method’s effectiveness and efficiency through thorough quantitative and qualitative experiments. Notably, it requires only about 12 hours of training on 8 A100 GPUs and 17M English data to perform very well on 110+ languages, with very close CLIP similarity scores of 38.61 (English) v.s. 37.61 (average of other languages).
@@ 2 Related Work
Recent advancements have seen the rise of diffusion-based models, which have significantly improved image generation quality and diversity. Popular models such as Stable Diffusion series [33, 11, 28], DALL-E [32, 31], Imagen [35] and Glide [24] demonstrate photorealistic generation capabilities, yet they are primarily trained on English data and thus struggle with multilingual image generation. Although diffusion models using CLIP text encoders can generalize somewhat to Romance languages (e.g., French), their performance significantly degrades for languages outside this family, particularly East Asian languages such as Chinese. This limitation arises from the models’ reliance on English-centric text encoders, such as CLIP [29] and T5 [30], and the predominantly English training data.
Recent efforts have explored multilingual image generation by incorporating multilingual text encoders and datasets to overcome the limitations of English-centric models. One approach involves building models entirely from scratch with non-English data. For instance, Hunyuan-DiT [19] and Kolors [40] incorporate Chinese text encoders and extensive Chinese datasets to enhance their support for culturally specific concepts and improve generation quality in Chinese. Alternatively, some methods attempt to adapt existing models by replacing or fine-tuning the text encoder. Models such as Taiyi [44], PanGu [22], and AltDiffusion [46] replace the text encoder in Stable Diffusion and then fine-tune it with multilingual data, thus reducing the overhead compared to training from scratch. Despite these efforts, these models still face computational challenges, as fine-tuning often requires updating the entire denoising diffusion UNet.
@@
Language models have evolved through various modeling approaches, reflecting a range of training objectives and capabilities. Early models, such as BERT [8], BART [17], focus on understanding sentence structure and contextual relationships by predicting masked tokens within a sentence. More recent large language models (LLMs), such as LLaMA series [42, 9], T5 [30] and GPT series [2, 25], build on this foundation with improved context understanding and generation abilities, allowing them to handle a wide range of tasks with nuanced language comprehension and open-ended text generation. Building on these foundational text encoders, recent works, such as CLIP [29], ALIGN [15], and InternVL [6], incorporate visual alignment through paired image-text data, creating joint representations that bridge language and vision for cross-modal tasks.
Previous works such as mBERT [8] and LLaMA3 [9] have shown strong multilingual capabilities by pre-training on large multilingual corpora, enabling these models to understand and generate text in a variety of languages. InternVL-LLaMA [6] achieves powerful multilingual cross-modal capabilities by aligning a text encoder with a vision transformer (ViT) on the multilingual image-text dataset LAION[36], enabling effective image-text contrastive learning. However, due to the scarcity of large-scale multilingual image-text datasets, some approaches [5, 3] have resorted to using translated datasets and distillation learning to align multilingual text features. Despite these advances, an open challenge remains inefficiently leveraging multilingual text encoders for text-to-image (T2I) generation. Previous works, such as Taiyi [44] and AltDiffusion [46], have adopted multilingual text encoders and fine-tuned them on multilingual image-text pairs for T2I tasks. In contrast, our method takes advantage of InternVL-LLaMA’s multilingual capabilities, requiring only a small amount of English data and without the need to fine-tune the SD model weights, achieving state-of-the-art performance in multilingual T2I generation.
@@ 3.1 Revisiting Text-to-Image Generation
BERT [8] encoder only MLM
T5 [30] encoder-decoder MLM
LLaMA [42, 9] decoder only NTP
CLIP-BERT [8] encoder only CL
InternVL-LLaMA [6] decoder-only NTP & CL
@@
Language Model. Existing T2I generation models typically utilize pre-trained language models as text encoders. For instance, the Stable Diffusion series [33, 28, 11] utilizes CLIP text encoder [29] or T5 [30] as language encoders. Beyond these encoder-only and encoder-decoder language models, recent years have seen increased attention on decoder-only large language models (LLMs), such as the GPT series [2, 25] and LLaMA [42, 9]. These models are trained on pure text data by predicting the next word, supporting multiple languages, and outperforming traditional language models in NLP tasks. As shown in Table 1, a wide variety of language models are now available in the community; however, determining which models are suited for multilingual text-to-image generation remains an open question.
@@
LAION-5B [36] 5B TI pairs 100+ Noisy
PixArt [4] 15M TI pairs 1 High
JourneyDB [26] 4M TI pairs 1 High
CCAligned [10] 100M Text Parallel 137 Noisy
CCMatrix [37] 69B Text Parallel 80+ Noisy
@@
Dataset. As shown in Table 2, the current datasets available in the community including multilingual text corpora, multilingual text-image paired datasets, and high-quality text-to-image datasets, hold potential value for multilingual image generation tasks. For example, large-scale datasets like LAION-400M/5B [36] contain large-scale noisy and lower-quality data, they provide multilingual text-image pairs, which are valuable for supporting multiple languages. Additionally, multilingual text translation corpora (such as CCMatrix [37]), while lacking corresponding image-text pair, offer useful cross-language correspondences. High-quality text-to-image datasets (such as JourneyDB [26]) provide high-resolution, high-quality images and are frequently used in text-image generation models; however, they primarily support English or some mainstream languages with limited coverage of other languages. Therefore, effectively utilizing low-cost, readily accessible internet-based text-image datasets and multilingual translation datasets to support multilingual text-to-image model training is a valuable area for further exploration.
@@ 3.2 MuLan: Toward Multilingual T2I Generation
Figure 2: Overview of of MuLan. The Transformer module is our Language Adapter, which follows the standard Transformer structure. We only train the Language Adapter, while all other modules are frozen. For different T2I model, we have designed different methods for using the Language Adapter.
(1) Language-Centered Alignment. A straightforward method that aligns multilingual semantics is to align multilingual semantics is to leverage a large set of translation data to align other languages’ vector spaces with the well-supported English vector space. By conducting distillation training with translation data alone, this can be achieved: we designate the Stable Diffusion language encoder as the teacher encoder, indicated by $L_{t}(x)$ and any multilingual encoder $L_{s}(x,\theta_{s})$ as the student encoder, aligning their features using MSE Loss. This alignment can be written as:
@@
Language Adapter. After getting the aligned language model, to achieve cost-effective multilingual text-to-image generation, we propose MuLan. This model incorporates a lightweight language adapter $L^{\prime}$ that bridges a multilingual-aligned language model with a visual generator, enabling generalization to multiple languages after training on a small amount of English text-to-image generation data $D_{\rm en}$. So the Eqn 1 could be rewritten as:
@@
In this work, we consider two types of adapters: MLP and transformer. These adapters are used to re-project the high-dimensional representations of text prompts from different languages into a unified space. We adopt different adapter designs for different diffusion models, as shown in Fig. 2. In detail, we can use either MLP or transfromer for projecting prompt embeddings for Pixart-$\alpha$ [4], and both architectures achieve good results. However, we find a simple MLP could not properly deal with Stable Diffusion models [33, 28]. Instead, we choose to use one layer encoder-decoder transformer with a set of learnable queries for extracting embeddings from InternVL outputs. For SDXL [28], we use two transformers to project embeddings for two text encoders that SDXL [28] adopts, and one attention pooling layer for extracting pool embeddings.
@@ 4 Experiments
This section describes the datasets, implementation details, and evaluation metrics used to assess our model, InternVL-MuLan. We compare its performance with state-of-the-art methods, including translation-based Stable Diffusion 1.5 [33] and AltDiffusion-m18 [46], across benchmarks such as XM18 [41], DPG-Bench [14], and COCO2014 [20]. Additionally, we perform ablation studies on adapter architectures and text encoders and present qualitative results to demonstrate our model’s capability to generate high-quality images across multiple languages.

### v2

@@ Abstract
In this work, we explore a cost-effective framework for multilingual image generation. We find that, unlike models tuned on high-quality images with multilingual annotations, leveraging text encoders pre-trained on widely available, noisy Internet image-text pairs significantly enhances data efficiency in text-to-image (T2I) generation across multiple languages. Based on this insight, we introduce MuLan, Multi-Language adapter, a lightweight language adapter with fewer than 20M parameters, trained alongside a frozen text encoder and image diffusion model. Compared to previous multilingual T2I models, this framework offers: (1) Cost efficiency. Using readily accessible English data and off-the-shelf multilingual text encoders minimizes the training cost; (2) High performance. Achieving comparable generation capabilities in over 110 languages with CLIP similarity scores nearly matching those in English (39.57 for English vs. 39.61 for other languages); and (3) Broad applicability. Seamlessly integrating with compatible community tools like LoRA, LCM, ControlNet, and IP-Adapter, expanding its potential use cases.
Machine Learning, ICML
@@ 1 Introduction
Recent diffusion models (Esser et al., 2024; Li et al., 2024a, b; Team, 2024; Wu et al., 2024; Zhang et al., 2022) for content generation have attained stunning advancements in terms of both aesthetic quality and text-content alignment. However, these models still face substantial limitations in multilingual support. For instance, one of the most popular image-generation models, Stable Diffusion (Rombach et al., 2022), and its successors (Podell et al., 2023; Esser et al., 2024) only supports English and a few Latin-based languages. This language barrier restricts the model’s performance in multilingual contexts and hinders its applicability worldwide across diverse cultural and linguistic backgrounds.
@@
The community primarily adopts two approaches to achieve multilingual text-to-image (T2I) generation. (1) The first is translation-based methods where input content is temporarily translated into English before generating the image. However, this approach often leads to inference delays, translation errors, and notable issues when handling slang or culturally nuanced content. (2) The other approach is native multilingual T2I models (Zhang et al., 2022; Radford et al., 2021; Arkhipkin et al., 2023; Shing & Akiba, ; Li et al., 2024b; Team, 2024), which is trained directly on high-quality images captioned in the target language. While this native approach improves image generation quality for non-English languages, it relies on extensive, carefully curated image-generation data in the target language, making it data- and resource-intensive. As a result, exploring a more efficient and generalizable approach to achieve strong multilingual generation capabilities remains challenging.
@@
On the other hand, thanks to the development of computational power and dataset scale, many existing language models have achieved strong multilingual capabilities through training on large-scale internet data. For example, models trained on text data (e.g., BERT (Devlin et al., 2019), the GPT series (Brown et al., 2020; OpenAI, 2023), and LLaMA (Touvron et al., 2023; Dubey et al., 2024)) and those trained on image-text pairs (e.g., CLIP (Radford et al., 2021), ALIGN (Jia et al., 2021), and InternVL (Chen et al., 2023b)) demonstrate outstanding performance in multilingual understanding. Since the multilingual capability of T2I generation models is closely tied to their text encoders, it becomes essential to explore how these powerful multilingual text encoders can be leveraged to enable existing generative models to achieve multilingual capabilities more efficiently and effectively.
@@
In this work, we deeply explore the application of multilingual semantic alignment in image generation from the perspective of language and image-text alignment. We also reveal that text encoders trained on large-scale multilingual image-text datasets with noisy data demonstrate remarkable data efficiency in multilingual image generation. Based on this insight, we introduce MuLan, a lightweight Multi-Language adapter that connects text encoders to diffusion models with low-cost training, facilitating native support for hundreds of languages. Specifically, we employ a plug-and-play language adapter with fewer than 20 million parameters to bridge a frozen multilingual text encoder with a frozen diffusion model, which exhibits outstanding zero-shot performance on multilingual T2I generation tasks.
@@
Our model offers advantages in terms of not only multilingual performance, but also training cost and adaptability compared to previous works. In a nutshell, MuLan leverages large-scale pre-trained multilingual text encoders, reducing the need for extensive multilingual datasets. As a result, the model can efficiently adapt to over 110 languages using only a small amount of English training data, achieving generation quality comparable to English (e.g., CLIP similarity scores: 39.57 for English, average 39.61 for other languages). Additionally, as shown in Figure 4, MuLan ’s plug-and-play architecture seamlessly integrates with existing model architectures and frameworks, greatly enhancing its compatibility with various community-driven tools and models, such as LoRA (Hu et al., 2021), LCM (Luo et al., 2023), ControlNet (Zhang et al., 2023), and IP-Adapter (Ye et al., 2023b).
@@
(1) We investigate different possible ways to equip monolingual text-to-image models with multilingual ability, among which we demonstrate that using a text encoder properly trained from noisy web-scale data (such as InternVL (Chen et al., 2023b)) is of great data efficiency.
@@
(3) We demonstrate our method’s effectiveness and efficiency through thorough quantitative and qualitative experiments. Notably, it requires only about 12 hours of training on 8 A100 GPUs and 17M English data to perform very well on 110+ languages, with very close CLIP similarity scores of 39.57 (English) v.s. 39.61 (average of other languages).
@@ 2 Related Work
Recent advancements have seen the rise of diffusion-based models, which have significantly improved image generation quality and diversity. Popular models such as SD series (Rombach et al., 2022; Esser et al., 2024; Podell et al., 2023), DALL-E (Ramesh et al., 2022, 2021), Imagen (Saharia et al., 2022) and Glide (Nichol et al., 2022) demonstrate photorealistic generation capabilities, yet they are primarily trained on English data and thus struggle with multilingual image generation. Although diffusion models using CLIP text encoders can generalize somewhat to Romance languages (e.g., French), their performance significantly degrades for languages outside this family, particularly East Asian languages such as Chinese. This limitation arises from the models’ reliance on English-centric text encoders, such as CLIP (Radford et al., 2021) and T5 (Raffel et al., 2020), and the predominantly English training data.
Recent efforts have explored multilingual image generation by incorporating multilingual text encoders and datasets to overcome the limitations of English-centric models. One approach involves building models entirely from scratch with non-English data. For instance, Hunyuan-DiT (Li et al., 2024b) and Kolors (Team, 2024) incorporate Chinese text encoders and extensive Chinese datasets to enhance their support for culturally specific concepts and improve generation quality in Chinese. Alternatively, some methods attempt to adapt existing models by replacing or fine-tuning the text encoder. Models such as Taiyi (Wu et al., 2024), PanGu (Lu et al., 2023), and AltDiffusion (Ye et al., 2023a) replace the text encoder in SD and then fine-tune it with multilingual data, thus reducing the overhead compared to training from scratch. GlueGen (Qin et al., 2023) and PEA-Diffusion (Ma et al., 2023) focus on fully utilizing the current T2I model by employing lightweight networks to align the semantics of other languages with English using low-cost parallel data. However, this often involves complex distillation loss, making the training process less stable. In our work, multilingual image generation capabilities can be generalized simply by training with the commonly used SD loss on English image-text pairs.
@@
Language models have evolved through various modeling approaches, reflecting a range of training objectives and capabilities. Early models, such as BERT (Devlin et al., 2019), BART (Lewis et al., 2019), focus on understanding sentence structure and contextual relationships by predicting masked tokens within a sentence. More recent large language models (LLMs), such as LLaMA series (Touvron et al., 2023; Dubey et al., 2024), T5 (Raffel et al., 2020) and GPT series (Brown et al., 2020; OpenAI, 2023), build on this foundation with improved context understanding and generation abilities, allowing them to handle a wide range of tasks with nuanced language comprehension and open-ended text generation. Building on these foundational text encoders, recent works, such as CLIP (Radford et al., 2021), ALIGN (Jia et al., 2021), and InternVL (Chen et al., 2023b), incorporate visual alignment through paired image-text data, creating joint representations that bridge language and vision for cross-modal tasks.
Previous works such as mBERT (Devlin et al., 2019) and LLaMA3 (Dubey et al., 2024) have shown strong multilingual capabilities by pre-training on large multilingual corpora, enabling these models to understand and generate text in a variety of languages. InternVL-LLaMA (Chen et al., 2023b) achieves powerful multilingual cross-modal capabilities by aligning a text encoder with a vision transformer (ViT) on the multilingual image-text dataset LAION(Schuhmann et al., 2022), enabling effective image-text contrastive learning. However, due to the scarcity of large-scale multilingual image-text datasets, some approaches (Chen et al., 2022; Carlsson et al., 2022) have resorted to using translated datasets and distillation learning to align multilingual text features. Despite these advances, an open challenge remains inefficiently leveraging multilingual text encoders for text-to-image (T2I) generation. Previous works, such as Taiyi (Wu et al., 2024) and AltDiffusion (Ye et al., 2023a), have adopted multilingual text encoders and fine-tuned them on multilingual image-text pairs for T2I tasks. In contrast, our method takes advantage of InternVL-LLaMA’s multilingual capabilities, requiring only a small amount of English data and without the need to fine-tune the SD model weights, achieving state-of-the-art performance in multilingual T2I generation.
@@ 3.1 Revisiting Text-to-Image Generation
BERT encoder only MLM
T5 encoder-decoder MLM
LLaMA decoder only NTP
CLIP encoder only CL
InternVL-LLaMA decoder-only NTP&CL
@@
Language Model. Existing T2I models typically employ pre-trained language models as text encoders. For instance, the SD series (Rombach et al., 2022; Podell et al., 2023; Esser et al., 2024) leverages CLIP (Radford et al., 2021) or T5 (Raffel et al., 2020). Recently, decoder-only LLMs like GPT (Brown et al., 2020; OpenAI, 2023) and LLaMA (Touvron et al., 2023; Dubey et al., 2024) have gained attention for their superior NLP performance, multilingual capabilities, and next-word prediction training. As summarized in Table 1, a diverse range of language models is available, yet the optimal choice for multilingual T2I generation remains an open question.
@@
LAION-5B 5B TI pairs 100+ Noisy
PixArt 15M TI pairs 1 High
JourneyDB 4M TI pairs 1 High
CCAligned 100M Text Parallel 137 Noisy
CCMatrix 69B Text Parallel 80+ Noisy
@@
Dataset. As shown in Table 2, existing datasets—including multilingual text corpora, multilingual text-image pairs, and high-quality text-to-image datasets—offer valuable resources for multilingual image generation. Large-scale datasets like Laion-400M/5B (Schuhmann et al., 2022) contain noisy, lower-quality data but provide multilingual text-image pairs, supporting diverse languages. Meanwhile, multilingual translation corpora (e.g., CCMatrix (Schwenk et al., 2020)) lack image-text pairs but offer cross-language alignments. High-quality text-to-image datasets (e.g., JourneyDB (Pan et al., 2023)) supply high-resolution images but mainly support English and a few mainstream languages. Thus, leveraging low-cost, internet-based text-image datasets and multilingual translation corpora to enhance multilingual T2I model training remains a promising direction.
@@ 3.2 MuLan: Toward Multilingual T2I Generation
Figure 2: Overview of of MuLan. We use a language model as the Multilingual Text Encoder, which has undergone Multilingual Semantic Alignment stage. We only train the Language Adapter, while all other modules are frozen.
(1) Language-Centered Alignment. A simple approach to multilingual semantic alignment is leveraging large-scale translation data to align other languages’ vector spaces with the well-supported English vector space. By conducting distillation training with translation data alone, this can be achieved: we designate the SD’s language encoder as the teacher encoder, indicated by $L_{t}(x)$ and any multilingual encoder $L_{s}(x,\theta_{s})$ as the student encoder, aligning their features using MSE Loss. This alignment can be written as:
@@
Language Adapter.
After getting the aligned language model, to achieve cost-effective multilingual text-to-image generation, we propose MuLan. This model incorporates a lightweight language adapter $L^{\prime}$ that bridges a multilingual-aligned language model with a visual generator, enabling generalization to multiple languages after training on a small amount of English text-to-image generation data $D_{\rm en}$. So the Eqn 1 could be rewritten as:
@@
These adapters are used to re-project the high-dimensional representations of text prompts from different languages into a unified space. We adopt different adapter designs for different diffusion models. In detail, we can achieve good results using a simple MLP architecture for Pixart-$\alpha$ (Chen et al., 2023a). However, we find MLP could not properly deal with SD models (Rombach et al., 2022; Podell et al., 2023). Instead, we choose to use one layer encoder-decoder transformer with a set of learnable queries for extracting embeddings from InternVL outputs. For SDXL (Podell et al., 2023), we use two transformers to project embeddings for two text encoders that SDXL adopts, and one attention pooling layer for extracting pool embeddings.
Bridge Together. As shown in Figure 2, after the Multilingual text encoder is trained with Multilingual Semantic Alignment, the Language Adapter can be trained with minimal English training data at a low cost while enabling multilingual image generation capabilities for broad text-to-image diffusion models. The Multilingual Semantic Alignment stage can be undergone with different strategies and base models, which could result in different performance on subsequent image-generation fine-tuning. More details are elucidated in Section 4.